# Uniform irreducibility of Galois action on the $\ell$-primary part of Abelian $3$-folds of Picard type

### v1

@@ Introduction
is not contained in a Borel subgroup. Manin’s original proof can be greatly simplified using Faltings’ proof of Mordell’s conjecture, which came later. In a series of papers Cadoret and Tamagawa established a definitive result regarding the uniform boundedness of the $\ell$-primary torsion for $1$-dimensional families of abelian varieties. In this paper we prove an analogous statement for certain $2$-dimensional families of abelian $3$-folds which we believe to be the first genuine result over a two-dimensional base.
Henceforth we fix an imaginary quadratic field $M$ of odd fundamental discriminant D-D and denote by $\mathcal{O}_{M}$ its ring of integers. An abelian $3$-fold of Picard type over a field $k$ containing $M$ will always stand for a principally polarized abelian variety over $k$ of dimension $3$ having multiplication by $\mathcal{O}_{M}$ defined over $k$. Its $\ell$-adic Tate module $T_{\ell}A$ is free of rank $3$ over $\mathcal{O}:=\mathbf{Z}_{\ell}\otimes\mathcal{O}_{M}$ endowed with a continuous $\mathcal{O}$-linear action of $\operatorname{Gal}_{k}$. By a line (resp. plane) in $T_{\ell}A$, we would mean a $\mathcal{O}$-submodule of rank $1$ (resp. $2$) which is a direct factor. More generally, given a positive integer $r$, a line (resp. plane) in $A[\ell^{r}]$ will always be assumed be the image, under the natural reduction map, of a line (resp. plane) in $T_{\ell}A$. Finally, by a full flag we would mean a tuple of a line sitting as direct factor in a plane. Lines (resp. planes) will be called $k$-rational if they are stable by $\operatorname{Gal}_{k}$ (but not necessarily point-wise fixed).
@@
We next show how one can relax the semi-stability assumption by adding a tiny bit of level structure at $D$. Given a prime $v$ of $M$ above some $p\mid D$, the projective $\operatorname{Gal}_{k}$-action on the $\mathbf{F}_{p}$-vector space $A[v]$ yields a homomorphism $\widetilde{\rho}_{A,p}:\operatorname{Gal}_{k}\to\operatorname{PGL}_{2}(\mathbf{F}_{p})$ (see (8)). Taking quotient by the unique index two subgroup $\operatorname{PSL}_{2}(\mathbf{F}_{p})$ of $\operatorname{PGL}_{2}(\mathbf{F}_{p})$ yields a canonical homomorphism $\varepsilon_{A,p}:\operatorname{Gal}_{k}\to\{\pm 1\}$ and we let $\varepsilon_{A,D}=\prod_{p\mid D}\varepsilon_{A,p}:\operatorname{Gal}_{k}\to\{\pm 1\}$.
@@
Theorem B is the main result of this paper and implies Theorem A as follows. Let $k^{\prime}$ be the compositum of the (finitely many) quadratic extensions of $k$ which are unramified outside $S$ and the primes dividing $D$. Given any abelian $3$-fold AA as in Theorem A, we claim that $\varepsilon_{A,D}(\operatorname{Gal}_{k^{\prime}})$ is trivial. Indeed, by a theorem of Grothendieck [14, Prop. 3.5] the semi-stability of $A$ at $v\notin S$, $v\nmid D$ implies that the inertia subgroup of $\operatorname{Gal}_{k}$ at $v$ acts unipotently on the $D$-adic Tate module of $A$, in particular its image by $\varepsilon_{A,D}$ is pro-$D$ hence trivial (as $D$ is odd). Therefore the base change of $A$ to $k^{\prime}$ satisfies the additional assumption in Theorem B, implying that Theorem A holds with $r(\ell,k^{\prime})$ from Theorem B.
@@
For an individual abelian variety $A$, the conclusion of Theorem B is a consequence of the Mumford–Tate conjecture which is known for abelian $3$-folds (see §2.3), so the important feature of the result is its uniformity. As abelian $3$-folds of Picard type are parametrized by Shimura surfaces of Picard type, a natural way to proceed would be to show that the $k$-rational points are not Zariski dense in any of their connected components $Y_{\Gamma}$. Let us for the moment consider the simpler situation from our earlier paper [10] where the congruence subgroups $\Gamma$ were neat. Our method there had two principal steps. The first step involved showing the existence of three linearly independent global holomorphic $1$-forms on the toroidal compactification $X_{\Gamma}$ (see §4.3 for an amended list of $\Gamma$ to which our methods apply). By a theorem of Faltings concerning the associated Albanese variety this implies that the $k$-rational points on $X_{\Gamma}$ are contained in a divisor $Z$, as predicted by a conjecture of Bombieri and Lang as $X_{\Gamma}$ turns out to be of general type. The second step consisted in applying a result of Nadel requiring $\Gamma$ to be neat and the canonical divisor to be big (in his sense) to deduce that any curve $C$ of genus $\leqslant 1$ contained in $X_{\Gamma}$ is in fact contained in the complement of $Y_{\Gamma}$. Consequently, every curve in $Z$ meeting the open surface $Y_{\Gamma}$ must be of genus $\geqslant 2$ thus, by Faltings’ proof of Mordell’s conjecture for curves, $Y_{\Gamma}(k)$ is finite for any number field $k$.
@@
Let us now say a few words about the techniques involved in the proof of Theorem B. As we are led to consider congruence subgroups of Iwahori type $\Gamma_{0}(\ell^{r})$, which are never neat as they have torsion, both steps mentioned above encounter difficulty and we have to resort to new methods. We produce irregularity by constructing explicit endoscopic automorphic forms in certain non-generic representations $\pi$ on the unitary group in $3$ variables. It is here that the index $2$ projective Galois image condition at $D$, suggested to us by Gross, is essential, as otherwise all of the Picard modular surfaces involved would have trivial Albanese and our approach would fail for global reasons. Making this strategy actually work yet requires to address some delicate representation theoretic questions to which a significant part of the paper is devoted and on which we will elaborate now.
@@
By Rogawski’s theory $\pi$ is an element of an endoscopic Arthur packet parametrized by an anti-cyclotomic (more precisely, conjugate-symplectic) Hecke character $\lambda$ of $M$. Theorem B imposes conditions so stringent so that $\lambda$ must differ from Gross’ minimally ramified ‘canonical’ characters by a finite order character $\chi$ only ramified at $\ell$. The local Arthur packet at $\ell$ contains two representations, a supercuspidal $\pi_{c,\ell}$ and a non-tempered $\pi_{n,\ell}$, both non-generic. The difficulty of finding $\Gamma_{0}(\ell^{r})$-invariants in $\pi_{c,\ell}$ forces us to work with the global $\pi_{n}$, which is automorphic if, and only if, the root number $W(\lambda^{3})$ is +1+1. For $D\equiv 3\pmod{8}$ Gross’ canonical characters work and a computation of matrix coefficients performed in §1.4 shows that the resulting $\pi_{n,\ell}$ has invariants even by the hyperspecial maximal compact subgroup.
@@
When $D\equiv 7\pmod{8}$ the canonical characters yield the wrong sign, leading us to consider $\lambda$’s which are tamely ramified at $\ell$ to switch the sign. It remains however to show that the non-tempered representations $\pi_{n,\ell}$ attached to such $\lambda$’s admit $\Gamma_{0}(\ell^{r})$-invariants for some $r$. For this we use a more involved argument, based on Jacquet modules and intertwining operators involving some precise averages of exponential sums, occupying the entirety section §1.5.
@@
Once the irregularity of $X_{\Gamma}$ has been shown to the at least $3$ and the Bombieri–Lang conjecture established, one has to deal with the possible curves $C$ of genus $\leqslant 1$ contained in $Y_{\Gamma}$. Using our key Lemma 4.1 affirming that our Picard modular surfaces only admit a finite number of isolated singularities, we show that, after removing a finite number of points, $C$ is endowed with an abelian family (see §2.2). This allows us apply the results of Cadoret and Tamagawa regarding the uniform boundedness of the Galois action on the Tate module of such $1$-dimensional families.
Finally, each of the finitely many non-CM $k$-rational points are dealt with using the results on the Mumford–Tate conjecture for abelian $3$-folds of Picard type recalled in §2.3.
@@
As our Picard modular surfaces $X_{\Gamma}$ have irregularity $q\geqslant 3$, the Kodaira–Spencer classification implies that they are either ruled of genus $q$, or elliptic, or else they are of general type. In the last case, which according to Holzapfel [16, §5.4] occurs for all odd $D\notin\{3,7,11,19,23,31,39,47,71\}$, we show that the Bombieri–Lang Conjecture holds, i.e., that the $k$-rational points are not Zariski dense. Investigating small values of $D$, as suggested by Mazur, seems even more interesting. It is established in loc. cit. that for all $D\neq 71$ in the above list the level $1$ Picard modular surfaces are rational and it would be natural to investigate the nature of their degree $2$ Gross covers that we consider. A way to shed light on this question would be to find an explicit $2$-parameter family of abelian $3$-folds of Picard type to which our theorem applies.
@@
The first author acknowledges support from the I-SITE ULNE grant ANR-16-IDEX-0004. The second author thanks the Simons Foundation for the grant 523557.
1. Levels for endoscopic non-tempered representations of $\mathrm{U}(3)$
It goes back to the work of Casselman that admissible irreducible representations having non-zero Iwahori invariants are exactly those occurring as sub-quotients in parabolic inductions of unramified characters. Whereas the dimension of the invariants by the depth $r$ Iwahori subgroup in the full induced representation grows as $r$ goes to infinity, this might not always be the case for all its sub-quotients, as shown by the example of the trivial representation of $\operatorname{GL}_{2}$, realized as a quotient of a unramified principal series representation.
@@ 1. Levels for endoscopic non-tempered representations of $\mathrm{U}(3)$
In this section we fully answer those two natural questions in the case of certain non-tempered endoscopic representations of $\operatorname{U}_{3}$ attached to a quadratic extension $E/\mathbf{Q}_{p}$. It will be later applied in a global setting to $E=M_{p}$, where $M$ is an imaginary quadratic field in which the prime $p$ does not split.
@@
Let $E$ be a quadratic field extension of $\mathbf{Q}_{p}$, $\mathcal{O}$ be its ring of integers, $\mathcal{P}$ its maximal ideal and $\varpi$ a uniformizer. We assume that $E$ is not a ramified extension of $\mathbf{Q}_{2}$. Denote by $x\mapsto\bar{x}$ the automorphism of $E$ induced by the non-trivial element of $\operatorname{Gal}(E/\mathbf{Q}_{p})$ and fix a generator $\xi$ of the different $\mathcal{D}$ of $E/\mathbf{Q}_{p}$ such that $\overline{\xi}=-\xi$. We fix an additive character $\psi:\mathbf{Q}_{p}\to\mathbf{C}^{\times}$ of conductor 0, i.e. $\ker(\psi)=\mathbf{Z}_{p}$, and we consider the additive character $\psi_{E}$ of $E$ defined as $\psi_{E}(x)=\psi(\operatorname{Tr}_{E/\mathbf{Q}_{p}}(x))$.
@@
$\langle x,y\rangle=\bar{x}_{1}y_{3}+\xi\bar{x}_{2}y_{2}-\bar{x}_{3}y_{1}.$
1.1. The Bruhat-Tits tree of $\operatorname{U}_{3}$
@@
As $G$ has rank $1$, its Bruhat-Tits building is a tree. We will first describe its standard apartment. The relative roots of $G$ are obtained by decomposing the adjoint action on the Lie algebra of the maximal $\mathbf{Q}_{p}$-split torus $T_{0}=\left\{\mathrm{diag}(a,1,a^{-1})|a\in\mathbf{Q}_{p}^{\times}\right\}$ of $G$. The positive elements of the associated root system $\Phi$ are $\{\zeta,2\zeta\}$. Let $h:\mathbf{G}_{m}\to T_{0}\subset G$ be the generator of the co-character lattice $X_{\ast}(T_{0})\simeq\mathbf{Z}$ such that $\langle\zeta,h\rangle=1$. Then the co-root sub-lattice is generated by $\zeta^{\vee}=2h$, so that we have the standard normalization $\langle\zeta,\zeta^{\vee}\rangle=2$. According to [32, §1.15] the affine roots are $\{\pm\zeta+\mathbf{Z}\}\cup\{\pm 2\zeta+\mathbf{Z}\}$ if $E$ is unramified, and $\{\pm\zeta+\tfrac{1}{2}\mathbf{Z}\}\cup\{\pm 2\zeta+\mathbf{Z}+\tfrac{1}{2}\}$ if $E$ is ramified; note that $\delta=0$ in loc. cit. as $E$ is not a ramified extension of $\mathbf{Q}_{2}$. The apartment associated to $T_{0}$ is $\mathbf{R}h$ and its walls are the vanishing sets of these (affine) roots, hence they are given by $\tfrac{1}{2}\mathbf{Z}h=\mathbf{Z}h\cup\tfrac{1}{2}\mathbf{Z}h$, resp. $\tfrac{1}{4}\mathbf{Z}h=\left(\tfrac{1}{2}\mathbf{Z}+\tfrac{1}{4}\right)h\cup\tfrac{1}{2}\mathbf{Z}h$, if $E$ is unramified, resp. ramified. Given an $\mathcal{O}$-lattice $\mathcal{L}$ in $E^{3}$ we define its dual as
$\mathcal{L}^{\perp}=\operatorname{Hom}_{\mathcal{O}}(\mathcal{L},\mathcal{D}^{-1})=\{x\in E^{3}|\langle x,\mathcal{L}\rangle\subset\mathcal{D}^{-1}\}.$
Lemma 1.1.
@@
There are two conjugacy classes of maximal compact subgroups in $G$, those which are stabilizers of self-dual lattices, and those which are stabilizers of almost self-dual lattices, i.e., lattices $\mathcal{L}$ such that $\mathcal{L}\subsetneq\mathcal{L}^{\perp}\subsetneq\varpi^{-1}\mathcal{L}$. They are all special, and the hyperspecial ones are those stabilizing a self-dual lattice when $E$ is unramified.
Proof.
A conjugacy class of maximal compact subgroups can be represented by a wall in the standard apartment. By definition, a wall is hyperspecial if for every $\zeta^{\prime}\in\Phi$ here exists an affine root with gradient $\zeta^{\prime}$ vanishing on that wall. Since $\left(\tfrac{1}{2}\mathbf{Z}+\tfrac{1}{4}\right)h\cap\tfrac{1}{2}\mathbf{Z}h=\varnothing$ this only can happen when $E$ is unramified, in which case the hyperspecial walls are $\mathbf{Z}h\cap\tfrac{1}{2}\mathbf{Z}h=\mathbf{Z}h$. All walls are special, as elements of $\Phi$ are rational multiple of one another. ∎
We now give an explicit description of the maximal compact subgroups corresponding to the walls of a chamber in the standard apartment.
@@
The standard maximal compact subgroup $K^{\circ}$ of $G$ is defined as the stabilizer of the self-dual lattice $\mathcal{L}^{\circ}=\mathcal{O}\oplus\xi^{-1}\mathcal{O}\oplus\xi^{-1}\mathcal{O}$. It is hyperspecial if and only if $E$ is unramified, and
$K^{\circ}=\begin{pmatrix}\mathcal{O}&\xi\mathcal{O}&\xi\mathcal{O}\\ \xi^{-1}\mathcal{O}&\mathcal{O}&\mathcal{O}\\ \xi^{-1}\mathcal{O}&\mathcal{O}&\mathcal{O}\end{pmatrix}\cap G.$
The reductive quotient $\overline{G}{}^{\circ}$ is given by $\operatorname{U}_{3}(\mathbf{F}_{p})$ if $E$ is unramified, and by $\mathrm{O}_{3}(\mathbf{F}_{p})$ if $E$ is ramified.
@@
The other standard maximal compact subgroup $K^{\prime}$ of $G$, defined as the stabilizer of the almost self-dual lattice $\mathcal{L}^{\prime}=\mathcal{O}\oplus\xi^{-1}\mathcal{O}\oplus\xi^{-1}\mathcal{P}$, is given by
$K^{\prime}=\begin{pmatrix}\mathcal{O}&\xi\mathcal{O}&\xi\mathcal{P}^{-1}\\ \xi^{-1}\mathcal{P}&\mathcal{O}^{\times}&\mathcal{O}\\ \xi^{-1}\mathcal{P}&\mathcal{P}&\mathcal{O}\end{pmatrix}\cap G.$
@@
One has $\mathcal{L}^{\prime\perp}=\mathcal{P}^{-1}\oplus\xi^{-1}\mathcal{O}\oplus\xi^{-1}\mathcal{O}$ and $K^{\prime}$ acts on $\varpi^{-1}\mathcal{L}^{\prime}/\mathcal{L}^{\prime\perp}\simeq\mathcal{O}/\mathcal{P}$ via its middle coefficient. The reductive quotient $\overline{G}{}^{\prime}$ is isomorphic to $\operatorname{U}_{1,1}(\mathbf{F}_{p})\times\operatorname{U}_{1}(\mathbf{F}_{p})$ if $E$ is unramified, and to $\pm\operatorname{SL}_{2}(\mathbf{F}_{p})\times\{\pm 1\}$, if $E$ is ramified.

### v2

@@ Abstract.
The first author acknowledges support from the I-SITE ULNE grant ANR-16-IDEX-0004. The second author thanks the Simons Foundation for support from the grant 523557.
@@ Introduction
is not contained in a Borel subgroup. Manin’s original proof can be greatly simplified using Faltings’ proof of Mordell’s conjecture, which came later. In a series of papers Cadoret and Tamagawa established a definitive result regarding the uniform boundedness of the $\ell$-primary torsion for $1$-dimensional families of abelian varieties. In this paper we prove an analogous statement for certain $2$-dimensional families of abelian $3$-folds which we believe to be the first result over a genuine two-dimensional base.
Henceforth we fix an imaginary quadratic field $M$ of odd fundamental discriminant $D\neq-3$ and denote by $\mathcal{O}_{M}$ its ring of integers. An abelian $3$-fold of Picard type over a field $k$ containing $M$ will always stand for a principally polarized abelian variety over $k$ of dimension $3$ having multiplication by $\mathcal{O}_{M}$ defined over $k$. Its $\ell$-adic Tate module $T_{\ell}A$ is free of rank $3$ over $\mathcal{O}:=\mathbf{Z}_{\ell}\otimes\mathcal{O}_{M}$ endowed with a continuous $\mathcal{O}$-linear action of $\operatorname{Gal}_{k}$. By a line (resp. plane) in $T_{\ell}A$, we would mean a $\mathcal{O}$-submodule of rank $1$ (resp. $2$) which is a direct factor. More generally, given a positive integer $r$, a line (resp. plane) in $A[\ell^{r}]$ will always be assumed be the image, under the natural reduction map, of a line (resp. plane) in $T_{\ell}A$. Finally, by a full flag we would mean a tuple of a line sitting as direct factor in a plane. Lines (resp. planes) will be called $k$-rational if they are stable by $\operatorname{Gal}_{k}$ (but not necessarily point-wise fixed).
@@
We next show how one can relax the semi-stability assumption by adding a tiny bit of level structure at $D$. Given a prime $v$ of $M$ above some $p\mid D$, the projective $\operatorname{Gal}_{k}$-action on the $\mathbf{F}_{p}$-vector space $A[v]$ yields a homomorphism $\widetilde{\rho}_{A,p}:\operatorname{Gal}_{k}\to\operatorname{PGL}(2,\mathbf{F}_{p})$ (see (15)). Taking quotient by the unique index two subgroup $\operatorname{PSL}(2,\mathbf{F}_{p})$ of $\operatorname{PGL}(2,\mathbf{F}_{p})$ yields a canonical homomorphism $\varepsilon_{A,p}:\operatorname{Gal}_{k}\to\{\pm 1\}$ and we let $\varepsilon_{A,D}=\prod_{p\mid D}\varepsilon_{A,p}:\operatorname{Gal}_{k}\to\{\pm 1\}$.
@@
Theorem B is the main result of this paper and implies Theorem A as follows. Let $k^{\prime}$ be the compositum of the (finitely many) quadratic extensions of $k$ which are unramified outside $S$ and the primes dividing $D$. Given any abelian $3$-fold AA as in Theorem A, we claim that $\varepsilon_{A,D}(\operatorname{Gal}_{k^{\prime}})$ is trivial. Indeed, by a theorem of Grothendieck [15, Prop. 3.5] the semi-stability of $A$ at $v\notin S$, $v\nmid D$ implies that the inertia subgroup of $\operatorname{Gal}_{k}$ at $v$ acts unipotently on the $D$-adic Tate module of $A$, in particular its image by $\varepsilon_{A,D}$ is pro-$D$ hence trivial (as $D$ is odd). Therefore the base change of $A$ to $k^{\prime}$ satisfies the additional assumption in Theorem B, implying that Theorem A holds with $r(\ell,k^{\prime})$ from Theorem B.
@@
For an individual abelian variety $A$, the conclusion of Theorem B is a consequence of the Mumford–Tate conjecture which is known for abelian $3$-folds (see §2.3), so the important feature of the result is its uniformity. As abelian $3$-folds of Picard type are parametrized by Shimura surfaces of Picard type, a natural way to proceed would be to show that the $k$-rational points are not Zariski dense in any of their connected components $Y_{\Gamma}$. Let us for the moment consider the simpler situation from our earlier paper [11] where the congruence subgroups $\Gamma$ were neat. Our method there had two principal steps. The first step involved showing the existence of three linearly independent global holomorphic $1$-forms on the toroidal compactification $X_{\Gamma}$. By a theorem of Faltings concerning the associated Albanese variety this implies that the $k$-rational points on $X_{\Gamma}$ are contained in a divisor $Z$, as predicted by a conjecture of Bombieri and Lang as $X_{\Gamma}$ turns out to be of general type. The second step consisted in applying a result of Nadel requiring $\Gamma$ to be neat and the canonical divisor to be big (in his sense) to deduce that any curve $C$ of genus $\leqslant 1$ contained in $X_{\Gamma}$ is in fact contained in the complement of $Y_{\Gamma}$. Consequently, every curve in $Z$ meeting the open surface $Y_{\Gamma}$ must be of genus $\geqslant 2$ thus, by Faltings’ proof of Mordell’s conjecture for curves, $Y_{\Gamma}(k)$ is finite for any number field $k$.
@@
Let us now say a few words about the techniques involved in the proof of Theorem B. As we are led to consider congruence subgroups of Iwahori type $\Gamma_{0}(\ell^{r})$, which are never neat as they have torsion, both steps mentioned above encounter difficulty and we have to resort to new methods. We produce irregularity by constructing explicit endoscopic automorphic forms in certain non-generic representations $\pi$ on the unitary group in $3$ variables. It is here that the index $2$ projective Galois image condition at $D$, suggested to us by Gross, is essential, as otherwise all of the Picard modular surfaces involved would have trivial Albanese and our approach would not apply for global reasons. Making this strategy actually work yet requires to address some delicate representation theoretic questions to which a significant part of the paper is devoted and on which we will elaborate now.
@@
By Rogawski’s theory $\pi$ is an element of an endoscopic Arthur packet parametrized by an anti-cyclotomic (more precisely, conjugate-symplectic) Hecke character $\lambda$ of $M$. Theorem B imposes conditions implying that $\lambda$ must differ from Gross’ minimally ramified ‘canonical’ characters by a finite order character of $M^{1}$ only ramified at $\ell$. The local Arthur packet at $\ell$ contains two representations, a supercuspidal $\pi_{c,\ell}$ and a non-tempered $\pi_{n,\ell}$, both non-generic. The difficulty of finding $\Gamma_{0}(\ell^{r})$-invariants in $\pi_{c,\ell}$ forces us to work with the global $\pi_{n}$, which is automorphic if, and only if, the global root number $W(\lambda^{3})$ equals +1+1. For $D\equiv 3\pmod{8}$ Gross’ canonical characters work and a computation of matrix coefficients performed in §1.4 shows that the resulting $\pi_{n,\ell}$ has invariants even by the hyperspecial maximal compact subgroup.
@@
When $D\equiv 7\pmod{8}$ the canonical characters yield the wrong sign, leading us to consider $\lambda$’s which are tamely ramified at $\ell$ to switch the sign. It remains however to show that the non-tempered representations $\pi_{n,\ell}$ attached to such $\lambda$’s admit $\Gamma_{0}(\ell^{r})$-invariants for some $r$, which is significantly harder than $\Gamma_{1}(\ell^{r})$-invariants. To that end, in §1.5 we devise a more involved argument relying on Jacquet modules and intertwining operators, and which requires precise computations of exponential sums beyond the reach of known estimates. It would be worthwhile exploring questions regarding levels of non-tempered representations in greater generality.
@@
Once the irregularity of $X_{\Gamma}$ has been shown to be at least $3$ and the Bombieri–Lang conjecture established, one has to deal with the possible curves $C$ of genus $\leqslant 1$ contained in $Y_{\Gamma}$. Using the key Lemma 2.1 on the lack of complex reflections, our Picard modular surfaces only admit a finite number of isolated singularities hence, after removing a finite number of points, $C$ is endowed with an abelian family (see §2.2). This allows us apply the results of Cadoret and Tamagawa regarding the uniform boundedness of the Galois action on the Tate module of such $1$-dimensional families. Finally, each of the finitely many non-CM $k$-rational points are dealt with using the Mumford–Tate conjecture for abelian $3$-folds of Picard type recalled in §2.3, and which is not available in higher dimension.
@@
As our Picard modular surfaces $X_{\Gamma}$ have irregularity $\geqslant 3$, the Enriques–Kodaira classification implies that they are either ruled of genus $q$, or elliptic, or else they are of general type. In the last case, which according to Holzapfel [17, §5.4] occurs for all odd $D\notin\{3,7,11,19,23,31,39,47,71\}$, we show that the Bombieri–Lang Conjecture holds, i.e., that the $k$-rational points are not Zariski dense. Investigating small values of $D$, as suggested by Mazur, seems even more interesting. It is established in loc. cit. that for all $D\neq 71$ in the above list the level $1$ Picard modular surfaces are rational and it would be natural to investigate the nature of their degree $2$ Gross covers that we consider. A way to shed light on this question would be to find an explicit $2$-parameter family of abelian $3$-folds of Picard type to which our theorem applies.
@@
1. Levels for endoscopic non-tempered representations of $\operatorname{U}(3)$
It goes back to the work of Casselman that admissible irreducible representations having non-zero Iwahori invariants are exactly those occurring as sub-quotients in parabolic inductions of unramified characters. Whereas the dimension of the invariants by the depth $r$ Iwahori subgroup in the full induced representation grows as $r$ goes to infinity, this might not always be the case for all its sub-quotients, as shown by the example of the trivial representation of $\operatorname{GL}(2)$, realized as a quotient of a unramified principal series representation.
@@ 1. Levels for endoscopic non-tempered representations of $\operatorname{U}(3)$
In this section we fully answer those two natural questions in the case of certain non-tempered endoscopic representations of $\operatorname{U}(3)$ attached to a quadratic extension $E/\mathbf{Q}_{p}$. It will be later applied in a global setting to $E=M_{p}$, where $M$ is an imaginary quadratic field in which the prime $p$ does not split.
@@
Let $E$ be a quadratic field extension of $\mathbf{Q}_{p}$, $\mathcal{O}$ be its ring of integers, $\mathcal{P}$ its maximal ideal and $\varpi$ a uniformizer. We assume that $E$ is not a ramified extension of $\mathbf{Q}_{2}$ and we fix a generator $\xi$ of its different ideal such that $\bar{\xi}=-\xi$. Denote by $x\mapsto\bar{x}$ the non-trivial element of $\operatorname{Gal}(E/\mathbf{Q}_{p})$. We fix an additive character $\psi:\mathbf{Q}_{p}\to\mathbf{C}^{\times}$ of conductor 0, i.e. $\ker(\psi)=\mathbf{Z}_{p}$, and we consider the additive character $\psi_{E}$ of $E$ defined as $\psi_{E}(z)=\psi(\operatorname{Tr}_{E/\mathbf{Q}_{p}}(z))$.
@@
$\langle x,y\rangle=\bar{x}_{1}y_{3}+\bar{x}_{2}y_{2}+\bar{x}_{3}y_{1}.$
The standard Borel $B$ of $G$ is a product of its torus
$T=\left\{\left(\begin{smallmatrix}\bar{\alpha}&&\\ &\beta&\\ &&\alpha^{-1}\end{smallmatrix}\right)\Big|\alpha\in E^{\times},\beta\in E^{1}\right\}$
with its unipotent subgroup
$N=\left\{[z,x]=\left(\begin{smallmatrix}1&-\bar{z}&\xi x-z\bar{z}/2\\ &1&z\\ &&1\end{smallmatrix}\right)\Big|z\in E,x\in\mathbf{Q}_{p}\right\}.$
1.1. The Bruhat-Tits tree of $\operatorname{U}(3)$
@@
As $G$ has rank $1$, its Bruhat-Tits building is a tree. We will first describe its standard apartment. The relative roots of $G$ are obtained by decomposing the adjoint action on the Lie algebra of the maximal $\mathbf{Q}_{p}$-split torus $T_{0}=\left\{\mathrm{diag}(a,1,a^{-1})|a\in\mathbf{Q}_{p}^{\times}\right\}$ of $G$. The positive elements of the associated root system $\Phi$ are $\{\zeta,2\zeta\}$. Let $h:\mathbf{G}_{m}\to T_{0}\subset G$ be the generator of the co-character lattice $X_{\ast}(T_{0})\simeq\mathbf{Z}$ such that $\langle\zeta,h\rangle=1$. Then the co-root sub-lattice is generated by $\zeta^{\vee}=2h$, so that we have the standard normalization $\langle\zeta,\zeta^{\vee}\rangle=2$. According to [32, §1.15] the affine roots are $\{\pm\zeta+\mathbf{Z}\}\cup\{\pm 2\zeta+\mathbf{Z}\}$ if $E$ is unramified, and $\{\pm\zeta+\tfrac{1}{2}\mathbf{Z}\}\cup\{\pm 2\zeta+\mathbf{Z}+\tfrac{1}{2}\}$ if $E$ is ramified; note that $\delta=0$ in loc. cit. as $E$ is not a ramified extension of $\mathbf{Q}_{2}$. The apartment associated to $T_{0}$ is $\mathbf{R}h$ and its walls are the vanishing sets of these (affine) roots, hence they are given by $\tfrac{1}{2}\mathbf{Z}h=\mathbf{Z}h\cup\tfrac{1}{2}\mathbf{Z}h$, resp. $\tfrac{1}{4}\mathbf{Z}h=\left(\tfrac{1}{2}\mathbf{Z}+\tfrac{1}{4}\right)h\cup\tfrac{1}{2}\mathbf{Z}h$, if $E$ is unramified, resp. ramified. A conjugacy class of maximal compact subgroups can be represented by a wall in the standard apartment. By definition, a wall is hyperspecial if for every $\zeta^{\prime}\in\Phi$ there exists an affine root with gradient $\zeta^{\prime}$ vanishing on that wall. Since $\left(\tfrac{1}{2}\mathbf{Z}+\tfrac{1}{4}\right)h\cap\tfrac{1}{2}\mathbf{Z}h=\varnothing$ this only can happen when $E$ is unramified, in which case the hyperspecial walls are $\mathbf{Z}h\cap\tfrac{1}{2}\mathbf{Z}h=\mathbf{Z}h$. All walls are special, as elements of $\Phi$ are rational multiple of one another.
Let us describe the conjugacy classes of maximal compact subgroups in $G$ in terms equivalent classes of $\mathcal{O}$-lattices $\mathcal{L}$ in $E^{3}$ modulo homothety. Recall the definition of the dual lattice
$\mathcal{L}^{\perp}=\operatorname{Hom}_{\mathcal{O}}(\mathcal{L},\mathcal{O})=\{x\in E^{3}|\langle x,\mathcal{L}\rangle\subset\mathcal{O}\}.$
@@
There are two conjugacy classes of maximal compact subgroups in $G$, those which are stabilizers of self-dual lattices, and those which are stabilizers of almost-self-dual lattices (i.e. $\mathcal{L}$ such that $\mathcal{L}\subsetneq\mathcal{L}^{\perp}\subsetneq\varpi^{-1}\mathcal{L}$). We next give an explicit description of the maximal compact subgroups corresponding to the walls of a chamber in the standard apartment.
@@
The standard maximal compact subgroup $K^{\circ}=G(\mathcal{O})$ of $G$ is defined as the stabilizer of the self-dual lattice $\mathcal{L}^{\circ}=\mathcal{O}^{3}$. It is hyperspecial if and only if $E$ is unramified. The reductive quotient $\overline{G}{}^{\circ}$ is given by $\operatorname{U}(3,\mathbf{F}_{p})$ if $E$ is unramified, and by $\mathrm{O}(3,\mathbf{F}_{p})$ if $E$ is ramified.
@@
The other standard maximal compact subgroup $K^{\prime}$ of $G$, defined as the stabilizer of the almost self-dual lattice $\mathcal{L}^{\prime}=\mathcal{O}\oplus\mathcal{O}\oplus\mathcal{P}$, is given by
$K^{\prime}=\left(\begin{smallmatrix}\mathcal{O}&\mathcal{O}&\mathcal{P}^{-1}\\ \mathcal{P}&\mathcal{O}^{\times}&\mathcal{O}\\ \mathcal{P}&\mathcal{P}&\mathcal{O}\end{smallmatrix}\right)\cap G.$
@@
One has $\mathcal{L}^{\prime\perp}=\mathcal{P}^{-1}\oplus\mathcal{O}\oplus\mathcal{O}$ and $K^{\prime}$ acts on $\mathcal{L}^{\prime}/\varpi\mathcal{L}^{\prime\perp}\simeq\mathcal{O}/\mathcal{P}$ via its middle coefficient. The reductive quotient $\overline{G}{}^{\prime}$ is isomorphic to $(\operatorname{U}(1,1)\times\operatorname{U}(1))(\mathbf{F}_{p})$ if $E$ is unramified, and to $\pm\mathbf{1}_{2}\cdot\operatorname{SL}(2,\mathbf{F}_{p})\times\{\pm 1\}$, if $E$ is ramified.